the service pool built a pool that was 100ft by 60ft wide. they want to built a sidewalk around the pool and had 450ft of fence. what is the maximum width that the sidewalk can be?
Answer by checkley77(12844)

2*100+2*60=200+120=320 is the perimeter of the pool.
2(100+2x)+2(60+2x)=450 this is the perimeter of the pool plus the sidewalk.
200+4x+120+4x=450
8x+320=450
8x=450-320
8x=130
x=130/8
x=16.25 ft is the width of the sidewalk.
proof:
2(100+32.5)+2(60+32.5)=450
2*132.5+2*92.5=450
265+185=450
450=450